OpenStudy (anonymous):

A recursive function is shown below: f(1) = 5 and f(n) = f(n - 1) - 6; n > 1 Which of the following lists the terms in the sequence defined by this recursive function? 5, 1, 7, 13, 19, ... 5, -1, -7, -13, -19, ... 5, 11, 17, 23, 29, ... 5, -11, -17, -23, -29, ...

OpenStudy (anonymous):

let's evaluate f(2), put n=2 in the relation:\[f(2)=f(2-1)-6=f(1)-6=5-6=-1\]

OpenStudy (anonymous):

would you calculate f(3)? give it a try and tell me what you get :-)